You decide to buy a house that costs 250,00. Your down payment is 20% of the price of the house and you take out a loan for the remaining balance. The bank charges an interest rate of 5% compounded quarterly. If you pay back the loan over 30 years what will be the amount of each monthly payment? please show the formula in your answer
Answer #1

Using financial calculator
Input:

Loan amount = PV = 250000*(1-0.2) = 200000

I/Y = 5/4 =1.25

N = 30*4 = 120

Solve for PMT as -3226.70

Hence the monthly payment = $3226.70